ON THE RANDOM CHARACTER OF FUNDAMENTAL CONSTANT EXPANSIONS

DAVID H. BAILEY AND RICHARD E. CRANDALL

Abstract. We propose a theory to explain random behavior for the digits in the expansions of fundamental mathematical constants. At the core of our approach is a general hypothesis concerning the distribution of the iterates generated by dynamical maps. On this main hypothesis, one obtains proofs of base-2 normality— namely bit randomness in a specific technical sense—for a collection of celebrated constants, including π, log 2, ζ(3), and others. Also on the hypothesis, the number ζ(5) is either rational or normal to base 2. We indicate a research connection between our dynamical model and the theory of pseudorandom number generators.